# UniTS: Unified Time Series Generative Model for Remote Sensing

### v1

@@ III-D Adaptive Condition Injector (ACor)
Figure 4: UniTS inference workflows for different time series tasks: (a) Multi-frame prediction for time series reconstruction and time series cloud removal, (b) Multi-frame prediction for time series semantic change detection, where class-specific maps are generated for each output frame, (c) Autoregressive multi-frame prediction for time series forecasting. Historical sequences and random noise are jointly fed into UniTS to predict the initial future frame. The predicted frames are then recursively used as conditions for subsequent time steps, progressively generating the full future sequence. To align with the temporal evolution characteristics of forecasting tasks, the input and spatio-temporal block are correspondingly adapted during both training and inference.
@@ III-E Spatiotemporal-aware Modulator (STM)
STM is designed to enhance the UniTS’s ability to capture complex spatio-temporal dependencies. By leveraging spatio-temporal prior information embedded in auxiliary data (e.g. Sentinel-1 unaffected by cloud cover), it generates a dynamic attention bias term that directly modulates the attention weights in the spatial/temporal attention module, thereby guiding UniTS to focus on regions with higher relevance in the spatio-temporal dimensions, as shown in Fig.3(c). Given auxiliary data ${\bf{q}}_{\scriptscriptstyle\text{STM}}$, we treat it as a dense, task-related prior signal (e.g., Sentinel-1 captures backscattering characteristics of different land cover types). STM encodes both the absolute positional relationship of feature map ${\bf{h}}_{\scriptscriptstyle\text{STM}}$ in flow matching and the latent relative geometric/evolutive relationship in the auxiliary data, constructing a learnable bias matrix $\mathbf{M}_{\text{spa}}\in\mathbb{R}^{n_{\text{h}}n_{\text{w}}\!\times\!n_{\text{h}}n_{\text{w}}}$ (for spatial attention) or $\mathbf{M}_{\text{tmp}}\in\mathbb{R}^{T\!\times\!T}$ (for temporal attention). This bias matrix is then injected into the attention score calculation as follows,
@@ III-F Training and Inference for Conditional Time Series Generation
  • In time series forecasting, the historical sequences of length $T_{\text{his}}$} are used as conditions, and the future sequences of length $T_{\text{fut}}$} serve as the target samples, with ${T_{\text{his}}}={T_{\text{fut}}}$ in this paper. While the above three tasks involve sampling the input noise into target samples for calculating Eq.(2), the forecasting task requires predicting both condition and target samples jointly during training phase, due to the complex temporal evolutionary relationship between the condition and target samples. This ensures the model effectively learns to predict future state changes while leveraging historical information. Furthermore, to prevent the temporal patterns of historical sequences from interfering with future predictions, we adjust the input and spatio-temporal block, as illustrated in Fig.4. The date list in metadata only includes future sequence dates to generate date embedding ${\bf{z}}_{\text{\tiny DOY}}^{\scriptscriptstyle\text{fut}}\in\mathbb{R}{{}^{{T_{\scriptscriptstyle\text{fut}}}\times d}}$. Additionally, a historical condition timestep embedding ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{con}}\in\mathbb{R}{{}^{{T_{\scriptscriptstyle\text{his}}}\times d}}$ is added to the flow time embedding. In the spatio-temporal block architecture, the ACor and STM components are removed from the temporal block. During training, ${\bf{z}}_{\text{\tiny DOY}}^{\scriptscriptstyle\text{fut}}$ is applied to the feature flow before the temporal block. After processing through multi-layer spatio-temporal blocks, the condition tokens ${\bf{z}}_{\text{con}}$ and ${\bf{\hat{z}}}_{t}^{\text{t}}$ are concatenated, while ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{con}}$ and ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{s}}$ are also concatenated and fed together into a linear decoder with AdaLN.
@@ IV-A Implementation Details
We follow the DiT architecture to implement UniTS. Unlike latent transformers that rely on pre-trained VAEs to compress video or time series inputs, UniTS directly tokenizes the pixel space. This is because most inputs to UniTS are multispectral images with significantly more channels than three, for which existing pre-trained VAE models are unsuitable. All experiments are trained using the AdamW optimizer with a constant learning rate of 1e41e-4. The input image size is fixed at $128\times 128$, and the Dopri5 solver is employed for sampling with 10 sample steps. The model, training hyperparameters, and inference settings are provided in Appendix A.
@@ IV-B Time Series Reconstruction
Figure 5: Qualitative comparison of time series reconstruction on TS-S12 Dataset, presenting the RGB band of Sentinel-2 here. The SSIM value of each frames in the time series is marked.
Figure 6: Qualitative comparison of time series cloud removal on TS-S12CR Dataset, presenting the RGB band of Sentinel-2 here. The SSIM value of each frames in the time series is marked. (Lavalleja Province, Uruguay, $(34^{\circ}29^{\prime}22.9^{\prime\prime}\mathrm{S},\ 55^{\circ}26^{\prime}24.0^{\prime\prime}\mathrm{W})$)
Qualitative Comparison. We select the area near Hipodromo in Tandil, Argentina and present in Fig.5 the time series reconstruction results of different methods for this region. Under clear or minimal cloud-mask coverage conditions, UniTS achieves over 98% SSIM relative to ground truth. For the frames heavily obscured by clouds or with severe data gaps (frames 1, 3-6, 8-11, 13, 16), methods such as VDPS, SiT, and VRT fail to reconstruct spatial details effectively, while U-TILISE produces noticeably blurred outputs across all reconstructed frames. UniTS maintains an average SSIM of 91% on missing frames and surpasses RESTORE-DiT and SeedVR in most frames, demonstrating superior capability in detail recovery and reconstruction of missing information.
Band-level and Class-level Reconstruction. We further analyze the reconstruction performance of UniTS under multi-modal conditional inputs for each Sentinel-2 spectral band and across different land cover classes. Fig.7(a) presents a heatmap of four metrics at the band-level, while Fig.7(b) displays the class-level overall performance score, obtained by normalizing and averaging five metrics. It can be observed from Fig.7(a) that the reconstruction performance of UniTS across wavelengths shows an increasing RMSE from the visible to the near-infrared range (B2-B8A), peaking at Red Edge 4 (B8A), and subsequently decreasing in the short-wave infrared bands (B11–B12). This trend is consistent with the results reported in RESTORE-DiT [8], and the underlying reasons are as follows:
  • Visible bands (B2-B4) show stable reflectance and high SNR, enabling the model to learn reliable mappings with minimal reconstruction error.
  • Red‑edge to NIR bands (B5-B8A), critical for vegetation response, are highly dynamic due to phenological changes, leading to degraded reconstruction. Band B8A, which is most sensitive to canopy structure and chlorophyll, is the hardest to reconstruct.
  • SWIR bands (B11-B12) are mainly affected by vegetation and soil moisture. Their slower variation and lower reflectance values constrain the error range, resulting in reduced reconstruction difficulty.
Among the nine land cover classes included in the TS-S12 dataset, Fig.7(b) shows that trees achieve the best reconstruction performance, while the four vegetation types, grass, flooded vegetation, crops, and shrub and scrub, exhibit comparable reconstruction scores, all maintained above 0.8. In contrast, performance drops markedly on snow and ice, whose high spectral variability (driven by snowpack conditions) and strong reflectance make them easily confused with clouds, raising reconstruction difficulty and limiting accurate spectral modeling.
@@ IV-C Time Series Cloud Removal
[…]
TABLE VII: Robustness of UniTS under modality missing. S1-Sentinel-1, S2-Real cloud-covered Sentinel-2.
Methods Training Inference PSNR↑ SSIM↑ RMSE↓ MAE↓ SAM↓
UniTS S2 S2 19.19 0.7391 0.1223 0.0890 7.83
UniTS S2 & S1 S2 19.44 0.7426 0.1187 0.0869 7.51
UniTS S2 & S1 S2 & S1 20.29 0.7592 0.1103 0.0828 7.42
@@
Qualitative Comparison. Fig.6 shows time series cloud removal over a suburban area in Lavalleja Province, Uruguay. The first three rows display Sentinel‑1, real cloud‑covered Sentinel‑2, and cloud‑free ground truth, with SSIM values given for each output. In this region, cloud coverage exceeds 90% across most 15‑day intervals; only the 9th and 10th frames capture partial ground information under thin clouds. All comparative methods perform poorly on heavily clouded frames. Even under such severe contamination, UniTS remains capable of recovering ground information closely resembling the cloud-free ground truth. Particularly in the last frame, leveraging its powerful spatiotemporal modeling and generative capabilities, UniTS obtains a cloud-free image with an SSIM of 0.9081.
Band-level and Class-level Cloud Removal. We further analyze the cloud removal performance of UniTS on severely cloud-contaminated Sentinel-2 imagery across different spectral bands and land cover classes. the band-level metric heatmap and the class-level overall performance score can be found in Appendix C. The experimental conclusions at the band-level and class-level are consistent with the observation results of the time series reconstruction task.
Robustness to Modality Absence During Inference. Multimodal cloud-removal methods often assume all modalities are available during inference. In practice, sensor failures or acquisition gaps (e.g., large-scale Sentinel-1 absences in TS-S12CR as shown in Fig.8) can leave certain modalities entirely missing. While approaches like U-TILISE suffer significant performance degradation when Sentinel-1 is absent or noisy, UniTS maintains stable cloud-removal quality even when Sentinel-1 is entirely absent. We quantitatively evaluate robustness under different training and inference modality configurations in Table VII. When trained on full modalities but tested with Sentinel-1 replaced by zero masks, UniTS shows only a moderate PSNR drop (-0.85 dB) and still substantially outperforms the single-modality (Sentinel-2 only) baseline. This indicates that:
  • UniTS learns flexible, modality‑balanced feature representations during training, avoiding over-dependence on any single source. This enables effective inference even under partial modality absence.
  • Multimodal training enriches the feature space and strengthens model generalization. Consequently, inference with missing modalities still yields higher performance than single-modality training.
The visualization showing the complete absence of Sentinel-1 during the inference process can be found in the Appendix D.
@@ IV-D Time Series Semantic Change Detection
TABLE VIII: Quantitative Comparison of Time Series Semantic Change Detection on DynamicEarthNet Dataset. The best and second-best values are bolded and underlined respectively.
@@
TABLE IX: Quantitative Comparison of Time Series Semantic Change Detection on MUDS Dataset. The best and second-best values are bolded and underlined respectively.
@@
Figure 9: Qualitative comparison of time series semantic change detection on DynamicEarthNet Dataset, presenting the RGB band of Planet here. Left: Semantic segmentation maps from T1 to T11, the mIoU value of each frames in the time series is marked; right: Binary Change Detection map (BCD) and Semantic Change Detection map (SCD).
Figure 10: Qualitative comparison of time series semantic change detection on MUDS Dataset. In the legend, TN, TP, FN, and FP respectively represent No Change, Change Detected, Missed Detection, and False Detection.
Dataset and Evaluation Settings. We evaluate our method on DynamicEarthNet [44] and MUDS [45]. The DynamicEarthNet dataset consists of Planet (RGBN) satellite time series images from ROIs worldwide collected between 2018 and 2019, comprising 54,750 images of size 1024×1024 with a spatial resolution of approximately 3 meters. It includes 24 monthly images annotated with land use and land cover classes. The MUDS dataset contains Planet (RGB) satellite time series images from 101 global ROIs acquired during 2018-2019, with monthly images annotated with building footprint labels. A sequence-to-sequence training and inference strategy is adopted, where both the conditional and target sample sequences are set to a a fixed length of $T=6$. The experiments compare against several state-of-the-art semantic change detection models (TSViT [37], U-TAE [13], A2Net [38], SCanNet [39] and TSSCD [14]) and multiple foundational remote sensing models (Scale-MAE [40], SatMAE++ [41], AnySat [42] and Skysense [43]). We employ multiple metrics to evaluate the performance of time series semantic change detection, including the mean intersection-over-union (mIoU), binary change score (BC), semantic change score (SC) and semantic change segmentation score (SCS).
Quantitative Comparison. Table VIII and Table IX present the per-class IoU and overall metrics of all methods on the two datasets. The strategy column in the table indicates the input-output configuration of each method: Mono denotes taking a single frame of time series as input and outputting a single segmentation map; Bi refers to taking two temporal images as input and producing a binary change mask along with two semantic segmentation maps; Multi represents taking the full time series as input and generating semantic segmentation results for each frame. UniTS demonstrates superior interpretation performance on both datasets compared to existing specific models and pre-trained foundational models, achieving mIoU scores of 42.52% and 61.96%, respectively. On one hand, this validates the advantage of the spatio-temporal block in UniTS in effectively capturing spatiotemporal dependencies and enabling high-precision temporal semantic segmentation. On the other hand, it shows that the representations learned through generative objectives such as flow matching can still exhibit strong generalization and spatiotemporal reasoning capabilities without relying on large-scale pre-training.
Qualitative Comparison. Fig.9 and Fig.10 respectively present the interpretation results of different methods on the two datasets. From the binary change detection maps in Fig.9, it can be observed that UniTS better detects changing regions in long time series and relatively accurately identifies transitions from soil to vegetation. In building change detection, UniTS maintains a detection accuracy of over 60% mIoU throughout the entire time series, with a low false detection rate in changing areas.
@@ IV-E Time Series Forecasting
TABLE X: Quantitative Comparison of Time Series Forecasting on TS-S12 Dataset. For full-band evaluation, the best and second-best values are bolded and underlined respectively. S2-Sentinel-2.
@@
TABLE XI: Quantitative Comparison of Time Series Forecasting on GreenEarthNet. The best and second-best values are bolded and underlined respectively.
@@
Figure 11: Qualitative comparison of time series forecasting on TS-S12 Dataset, presenting the RGB band of Sentinel-2 here. The SSIM value of each frames in the time series is marked. (Boston, USA, $(42^{\circ}18^{\prime}37.8^{\prime\prime}\mathrm{N},\ 71^{\circ}06^{\prime}31.0^{\prime\prime}\mathrm{W})$)
@@
Qualitative Comparison. Fig.11 shows time series forecasting over Franklin Park, Boston, USA. The first row presents historical sequences and future ground-truth sequences; subsequent rows display outputs from different methods. Unlike existing video-generation or forecasting approaches, UniTS not only produces spatially coherent, detail-rich future frames but also more accurately captures temporal climate and phenological evolution. Band-level reflectance time series (in Fig.11) further confirm UniTS’s strength in modeling complex spatiotemporal dynamics, ensuring both spatiotemporal consistency and phenological accuracy in its predictions.
@@ IV-F Ablation Study
TABLE XII: Ablation Comparison of Different Modules in UniTS.
@@
TABLE XIII: Ablation Comparison of Different Conditional Fusion.
@@
TABLE XIV: Ablation Comparison of Sampling Steps in Flow Matching.
@@
Effect of Metadata, Acor and STM. Table XII presents the ablation results of different modules in UniTS. The introduction of the ACor module brings significant performance improvement, particularly in the time series cloud removal task, where the PSNR increases by approximately 1.44 dB, demonstrating its effectiveness in dynamically fusing multimodal conditional information. The STM module further optimizes spatiotemporal dependency modeling, especially in forecasting tasks. The metadata provides essential spatiotemporal priors for the model, playing a notable role in enhancing forecasting accuracy. When all three modules are enabled simultaneously, the model achieves the optimal performance, indicating functional complementarity among the modules and their collective contributions to building comprehensive spatiotemporal understanding and generative capabilities.
Condition Fusion Strategy. We compare the impact of different conditional fusion strategies on the performance of UniTS in Table XIII. Simple feature concatenation and cross-attention mechanisms perform similarly in both time series cloud removal and forecasting tasks, indicating limitations of traditional fusion approaches in modeling spatiotemporal features. In contrast, the Acor-based fusion strategy achieves the best performance, demonstrating that Acor can more effectively coordinate multimodal conditional information by dynamically generating affine transformation parameters.
Flow Matching Samples Steps. In Table XIV, we evaluate the impact of the number of sampling steps in the flow matching on generation performance. As the number of sampling steps increases from 10 to 40, only minor fluctuations are observed in PSNR, SSIM, and SAM, with no clear monotonic upward or downward trend. This indicates that UniTS can achieve high-quality sequence generation with only a small number of sampling steps.
IV-G Limitations and Future Research
Limitations. Data: Although our constructed TS-S12 and TS-S12CR datasets are derived from Sentinel-1 & 2 data of over ten thousand ROIs worldwide, the field of remote sensing encompasses multi-source time series imagery with varying resolutions and spectral bands. Relying solely on these two datasets is insufficient for conducting multi-source data collaborative tasks such as time series cloud removal and forecasting. Methodology: The main limitation of UniTS lies in its direct learning of the mapping from noise to targets in pixel space without constructing a latent space. This restricts UniTS’s applicability across multiple resolutions, particularly when dealing with large-scale remote sensing imagery.
Future research. Data: We will extend platform diversity and geographic coverage by integrating multi‑source data (Planet, Landsat‑8/9, MODIS, etc). Methodology: 1) We will explore latent‑space representations tailored for remote sensing multimodal data, adopting advanced techniques such as Representation Autoencoders (RAE) [54] and SVG [55], and enhance generation quality with architectures like DiT‑3D. 2) In the context of the rapid development of world models like DINO-world, video/time series forecasting has become a core capability for building world models. UniTS, as a unified generative framework for remote sensing spatiotemporal data, has made a preliminary attempt in this direction. In the future, we will dedicate efforts to constructing a world model specifically tailored for the earth observation.

### v2

@@ III-D Adaptive Condition Injector (ACor)
Figure 4: UniTS inference workflows for different time series tasks: (a) Multi-frame prediction for time series reconstruction and time series cloud removal, (b) Multi-frame prediction for time series semantic change detection, where class-specific maps are generated for each output frame, (c) Autoregressive multi-frame prediction for time series forecasting. Historical sequences and random noise are jointly fed into UniTS to predict the initial future frame. The predicted frames are then recursively used as conditions for subsequent time steps, progressively generating the full future sequence. To align with the temporal evolution characteristics of forecasting tasks, the input and spatiotemporal block are correspondingly adapted during both training and inference.
@@ III-E Spatiotemporal-aware Modulator (STM)
STM is designed to enhance the UniTS’s ability to capture complex spatiotemporal dependencies. By leveraging spatiotemporal prior information embedded in auxiliary data (e.g. Sentinel-1 unaffected by cloud cover), it generates a dynamic attention bias term that directly modulates the attention weights in the spatial/temporal attention module, thereby guiding UniTS to focus on regions with higher relevance in the spatiotemporal dimensions, as shown in Fig.3(c). Given auxiliary data ${\bf{q}}_{\scriptscriptstyle\text{STM}}$, we treat it as a dense, task-related prior signal (e.g., Sentinel-1 captures backscattering characteristics of different land cover types). STM encodes both the absolute positional relationship of feature map ${\bf{h}}_{\scriptscriptstyle\text{STM}}$ in flow matching and the latent relative geometric/evolutive relationship in the auxiliary data, constructing a learnable bias matrix $\mathbf{M}_{\text{spa}}\in\mathbb{R}^{n_{\text{h}}n_{\text{w}}\!\times\!n_{\text{h}}n_{\text{w}}}$ (for spatial attention) or $\mathbf{M}_{\text{tmp}}\in\mathbb{R}^{T\!\times\!T}$ (for temporal attention). This bias matrix is then injected into the attention score calculation as follows,
@@ III-F Training and Inference for Conditional Time Series Generation
  • In time series forecasting, the model uses historical sequences of length $T_{\text{his}}$ as conditions to predict future sequences of length $T_{\text{fut}}$, with $T_{\text{his}}=T_{\text{fut}}$ in this paper. Unlike the previous tasks, forecasting requires jointly modeling both condition and target samples during training to capture their temporal dependencies. To prevent historical patterns from interfering with future predictions, we adjust the input and spatiotemporal block architecture, as shown in Fig. 4. Specifically, the date embedding ${\bf{z}}_{\text{\tiny DOY}}^{\scriptscriptstyle\text{fut}}\in\mathbb{R}{{}^{{T_{\scriptscriptstyle\text{fut}}}\times d}}$ is generated using only future dates from the metadata. A historical condition timestep embedding ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{con}}\in\mathbb{R}{{}^{{T_{\scriptscriptstyle\text{his}}}\times d}}$ is added to the flow time embedding. In the temporal block, the ACor and STM components are removed. During training, ${\bf{z}}_{\text{\tiny DOY}}^{\scriptscriptstyle\text{fut}}$ is applied to the feature flow before the temporal block. After passing through spatiotemporal blocks, the condition tokens $\mathbf{z}_{\text{con}}$ and $\hat{\mathbf{z}}_{t}^{\text{t}}$ are concatenated, along with ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{con}}$ and ${\bf{z}}_{\scriptscriptstyle\text{FM}}^{\text{s}}$, and fed into a linear decoder with AdaLN.
@@ IV-A Implementation Details
We follow the DiT architecture to implement UniTS. Unlike latent transformers that rely on pre-trained VAEs to compress video or time series inputs, UniTS directly tokenizes the pixel space. This is because most inputs to UniTS are multispectral images with significantly more channels than three, for which existing pre-trained VAE models are unsuitable. All experiments are trained using the AdamW optimizer with a constant learning rate of 1e41e-4. The input image size is fixed at $128\times 128$, and the Dopri5 solver is employed for sampling with 10 sample steps. The model, training hyperparameters, and inference settings are provided in Appendix A. The qualitative comparison of all experiments are provided in Appendix C-G.
@@ IV-B Time Series Reconstruction
[…]
@@
Band-level and Class-level Reconstruction. We further analyze UniTS’s reconstruction performance under multi-modal conditional inputs across Sentinel-2 spectral bands and land cover classes. Fig. 6(a) shows band-level metrics via heatmap, while Fig. 6(b) presents the class-level score averaged from five metrics. As shown in Fig. 6(a), UniTS exhibits increasing RMSE from visible to near-infrared bands (B2–B8A), peaking at Red Edge 4 (B8A), then decreasing in short-wave infrared bands (B11–B12), a trend consistent with RESTORE-DiT [3]. Among the nine land cover classes in the TS-S12 dataset, trees achieve the highest reconstruction performance (Fig. 6(b)). The four vegetation types, grass, flooded vegetation, crops, and shrub/scrub, show similar scores. In contrast, snow and ice exhibit significantly lower performance due to their high spectral variability and strong reflectance, which complicates their distinction from clouds and accurate spectral modeling.
@@ IV-C Time Series Cloud Removal
Band-level and Class-level Cloud Removal. We further analyze the cloud removal performance of UniTS on severely cloud-contaminated Sentinel-2 imagery across different spectral bands and land cover classes. the band-level metric heatmap and the class-level overall performance score can be found in Appendix D. The experimental conclusions at the band-level and class-level are consistent with the observation results of the time series reconstruction task.
Robustness to Modality Absence During Inference. Multimodal cloud-removal methods often assume all modalities are available during inference. However, sensor failures or acquisition gaps can leave certain modalities entirely missing. While approaches like U-TILISE suffer significant performance degradation when Sentinel-1 is absent or noisy, UniTS maintains stable cloud-removal quality even when Sentinel-1 is entirely absent. We evaluate robustness under different training and inference modality configurations in Fig.5. When trained on full modalities but tested with Sentinel-1 replaced by zero masks, UniTS shows only a moderate PSNR drop (-0.85 dB) and still outperforming the Sentinel-2-only baseline.
@@ IV-D Time Series Semantic Change Detection
TABLE VII: Quantitative Comparison of Time Series Semantic Change Detection on DynamicEarthNet Dataset. The best and second-best values are bolded and underlined respectively.
@@
TABLE VIII: Quantitative Comparison of Time Series Semantic Change Detection on MUDS Dataset. The best and second-best values are bolded and underlined respectively.
@@
Dataset and Evaluation Settings. We evaluate our method on DynamicEarthNet [38] and MUDS [39]. DynamicEarthNet contains monthly Planet satellite images from 2018–2019, with 54,750 RGBN images of size 1024×1024 and 3-meter resolution, annotated for land use and land cover. MUDS includes monthly Planet RGB images from 101 global regions in the same period, annotated with building footprints. A sequence-to-sequence training and inference strategy is adopted, where both the conditional and target sample sequences are set to a a fixed length of $T=6$. The experiments compare against several state-of-the-art semantic change detection models (TSViT [31], U-TAE [6], A2Net [32], SCanNet [33] and TSSCD [7]) and multiple foundational remote sensing models (Scale-MAE [34], SatMAE++ [35], AnySat [36] and Skysense [37]). We employ multiple metrics to evaluate the performance, including the mean intersection-over-union (mIoU), binary change score (BC), semantic change score (SC) and semantic change segmentation score (SCS).
Quantitative Comparison. Table VII and Table VIII present the per-class IoU and overall metrics of all methods on the two datasets. The strategy column indicates the input-output setup: Mono uses a single frame to produce one segmentation map; Bi takes two temporal images and outputs a binary change mask plus two semantic maps; Multi takes the full time series and generates segmentation for each frame. UniTS achieves the best interpretation performance on both datasets, with mIoU scores of 42.52% (DynamicEarthNet) and 61.96% (MUDS). These results highlight the effectiveness of UniTS’s spatiotemporal block in capturing spatiotemporal dependencies and performing precise temporal semantic segmentation. They also demonstrate that representations learned through flow matching generative objectives retain strong generalization and spatiotemporal reasoning ability without requiring large‑scale pretraining.
@@ IV-E Time Series Forecasting
TABLE IX: Quantitative Comparison of Time Series Forecasting on TS-S12 Dataset. For full-band evaluation, the best and second-best values are bolded and underlined respectively. S2-Sentinel-2.
@@
TABLE X: Quantitative Comparison of Time Series Forecasting on GreenEarthNet. The best and second-best values are bolded and underlined respectively.
@@ IV-F Ablation Study
TABLE XI: Ablation Comparison of Different Modules in UniTS.
@@
TABLE XII: Ablation Comparison of Different Conditional Fusion.
@@
TABLE XIII: Ablation Comparison of Sampling Steps in Flow Matching.
@@
Effect of Metadata, Acor and STM. Table XI presents the ablation results of different modules in UniTS. The introduction of the ACor module brings significant performance improvement, particularly in the time series cloud removal task, where the PSNR increases by approximately 1.44 dB, demonstrating its effectiveness in dynamically fusing multimodal conditional information. The STM module further optimizes spatiotemporal dependency modeling, especially in forecasting tasks. The metadata provides essential spatiotemporal priors for the model, playing a notable role in enhancing forecasting accuracy. When all three modules are enabled simultaneously, the model achieves the optimal performance, indicating functional complementarity among the modules and their collective contributions to building comprehensive spatiotemporal understanding and generative capabilities.
Condition Fusion Strategy. We compare the impact of different conditional fusion strategies on the performance of UniTS in Table XII. Simple feature concatenation and cross-attention mechanisms perform similarly in both time series cloud removal and forecasting tasks, indicating limitations of traditional fusion approaches in modeling spatiotemporal features. In contrast, the Acor-based fusion strategy achieves the best performance, demonstrating that Acor can more effectively coordinate multimodal conditional information by dynamically generating affine transformation parameters.
Flow Matching Samples Steps. In Table XIII, we evaluate the impact of the number of sampling steps in the flow matching on generation performance. As the number of sampling steps increases from 10 to 40, only minor fluctuations are observed in PSNR, SSIM, and SAM, with no clear monotonic upward or downward trend. This indicates that UniTS can achieve high-quality sequence generation with only a small number of sampling steps.